OpenStudy (anonymous):

Meal/ Serv. Good/ serv. poor/ total Lunch/ 32 / 15 / 47 Dinner/ 48 /32 / 80 Total / 80 / 47 /127 The table shows the results of a restaurant survey. Find the probability the service was poor, given that the meal was dinner. Help please!

OpenStudy (turingtest):

Let A = the service was poor Let B = the meal was dinner we are looking for the conditional probability of A given B, which is\[P(A|B)=\frac{P(A\cap B)}{P(B)}\] What is the probability that the meal was dinner and poor? What is the meal was dinner?

OpenStudy (anonymous):

Thank you so much! I figured it out. It's the number service poor has in common with dinner which is 32 divided by the total number of dinners 80, therefore, 32 /80= 2/5 :)